OpenStudy (anonymous):

Determine which of the following are the solutions to the equation below. x^2=11

OpenStudy (xapproachesinfinity):

what do you think

OpenStudy (xapproachesinfinity):

any attempt

OpenStudy (anonymous):

i have no clue how to do it

OpenStudy (xapproachesinfinity):

how do we solve \[x^2=1\]

OpenStudy (anonymous):

idk

OpenStudy (xapproachesinfinity):

ok let's ask a different way what number is i square it gives me one

OpenStudy (xapproachesinfinity):

so?

OpenStudy (xapproachesinfinity):

do we have \[1^2=1~~and~~(-1)^2=1\]

OpenStudy (xapproachesinfinity):

so two number result in 1

OpenStudy (anonymous):

so 11^2 = 22

OpenStudy (xapproachesinfinity):

no that is now my point

OpenStudy (xapproachesinfinity):

my point is some number or numbers if i square it i will get

OpenStudy (xapproachesinfinity):

when we have \[x^2=a\] to find what x is we take square root of both sides \[\sqrt{x^2}=\sqrt{a}\] square root cancels with the square then get only \[|x|=\sqrt{a}\] which is the same as \[x=\pm\sqrt{a}\]

OpenStudy (xapproachesinfinity):

in your example you have a=11 so just replace a with 11

OpenStudy (anonymous):

so the answer is 11

OpenStudy (xapproachesinfinity):

no! look carefully at what i did and do the same

OpenStudy (xapproachesinfinity):

take the square root of both sides first

OpenStudy (anonymous):

\[\sqrt{x^2} and \sqrt{11}\]

OpenStudy (xapproachesinfinity):

put equal sign \[\sqrt{x^2}=\sqrt{11}\]

OpenStudy (xapproachesinfinity):

go on...

OpenStudy (anonymous):

\[\sqrt{x^2}=\sqrt{11}\]

OpenStudy (xapproachesinfinity):

follow and do the rest

OpenStudy (anonymous):

\[\left| x \right|=\sqrt{11}\]

OpenStudy (xapproachesinfinity):

aha..

OpenStudy (anonymous):

\[x=\pm \sqrt{11}\]

OpenStudy (anonymous):

is that right

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

and thats all right

OpenStudy (xapproachesinfinity):

yes